OpenStudy (milo123):

MEDAL What is the volume of a sphere with a radius of 6 inches? 112π cubic inches 288π cubic inches 290π cubic inches 300π cubic inches

OpenStudy (milo123):

@Elsa213 @Will.H @mathmate

OpenStudy (will.h):

\[V = \frac{ 3 }{ 4 } * \pi * r^3\] \[\[V = \frac{ 3 }{ 4 } * \pi * 6^3\]\] V = 904.78

OpenStudy (milo123):

So it's D?

OpenStudy (will.h):

Actually B

OpenStudy (phi):

the equation is \[ V = \frac{4}{3} \pi \ r^3 \] if r is 6, replace the r with "6" \[ \frac{4}{3} \pi \ 6^3 \] or , changing the order of the multiplies, \[ \frac{4}{3} \cdot \ 6^3 \pi \]

OpenStudy (phi):

now you have to figure out 4*6*6*6/3 6/3 is 2, so we can change that to 4*6*6*2 2*4 is 8 , so that is also 8*6*6 6*8 is 48 48*6 finally 6*48 is 240+48 = 288 and don't forget the pi \[ V = 288 \pi \]

OpenStudy (milo123):

thanks